A ball falls from a height of 5m and strikes the roof of a lift. If at the time of collision lift is moving in upward direction with velocity of 1m/s then the velocity with which the ball rebounds after collision will be

Grade:11

3 Answers

Arun
5 years ago
Dear student
 
I am getting my answer as 12m/sec. Please lef me know if this is correct then I will share my solution
 
Hope it helps
 
Regards
Arun (askIITians forum expert)
Akshay keshwani
5 years ago
Please share the solution with explanation............................................................
Krutarth
5 years ago
MASS OF LIFT = M2                               MASS OF BALL = M1
V1=(M2-M1)​​         2M2  U2
       M2+M1            M2+M1
 
                                                M2>>M1
V1=M2  U1       +     2M2  U2
        M2                         M2
 
=U1+2U2
 
h=5m     ,    v2-u2=2gh
           0  -u12 =2x10x5...u1=10
             v1 = 10+2x1(u1+2u2)
                 =12m/s